# ARIM-mdx Data System: Towards a Nationwide Data Platform for Materials Science

### v1

@@ Abstract
In modern materials science, effective and high-volume data management across leading-edge experimental facilities and world-class supercomputers is indispensable for cutting-edge research. Such facilities and supercomputers are typically utilized by a wide range of researchers across different fields and organizations in academia and industry. However, existing integrated systems that handle data from these resources have primarily focused just on smaller-scale cross-institutional or single-domain operations. As a result, they often lack the scalability, efficiency, agility, and interdisciplinarity, needed for handling substantial volumes of data from various researchers.
In this paper, we introduce ARIM-mdx data system, a nationwide data platform for materials science in Japan. The platform involves 8 universities and institutes all over Japan through the governmental materials science project. Currently in its trial phase, the ARIM-mdx data system is utilized by over 800 researchers from around 140 organizations in academia and industry, being intended to gradually expand its reach. The system employs a hybrid architecture, combining a peta-scale dedicated storage system for security and stability with a high-performance academic cloud for efficiency and scalability. Through direct network connections between them, the system achieves 4.7x latency reduction compared to a conventional approach, resulting in near real-time interactive data analysis. It also utilizes specialized IoT devices for secure data transfer from equipment computers and connects to multiple supercomputers via an academic ultra-fast network, achieving 4x faster data transfer compared to the public Internet. The ARIM-mdx data system, as a pioneering nationwide data platform, has the potential to contribute to the creation of new research communities and accelerates innovations.
@@
Big Data, Materials Science, Supercomputer, Cloud Computing, Shared-use Research Facility, Governmental Project
@@ I Introduction
Figure 1: Collaboration between experimental and theoretical materials science. The example shows a typical nonostructure analysis, where the atomic-scale image of the targeting material is obtained from the electron microscope whereas it is analyzed via computation-intensive theoretical simulation using the supercomputer to detect its structure.
@@
However, despite the substantial success in the wide collection and large-scale management of the computational data, that of the real experimental data remains a significant challenge. Most efforts on the collection, management and analysis of the experimental data, partly integrated with the computational data, are primarily focused just on the smaller-scale cross-institution [12, 13, 14, 15, 16, 17, 18, 19], single domain [20, 21, 22, 23, 24, 25], or data curated from research articles [26, 27]. This leaves a huge potential for leveraging a large amount of the raw, unprocessed data untouched.
To address these gaps, we have been developing ARIM-mdx data system, a nationwide materials-science data infrastructure capable of effectively handling both computational and experimental data. As a foundational infrastructure to support modern research collaboration, the ARIM-mdx data system is designed to meet essential system requirements for data storage, management, and analysis. The system requirements and our contributions for this nationwide deployment are summarized as follows:
@@
For the stability and security, the ARIM-mdx data system is based on a peta-scale dedicated storage system, rather than on public clouds or shared-use storage systems. Such an isolated setup enables us to make physical operations, such as power managements or physical network configurations, be fully under control. Moreover, to ensure safe data transfer from network-disabled facility computers to the ARIM-mdx data system, we offload the data transfer mechanism to IoT devices, thereby mitigating security risks on the facility computers.
Scalability and Efficiency: Due to the rapid advancement of experimental facilities and supercomputers, these generate richer data at higher rate. For example, the state-of-the-art Cryogenic Electron Microscopy (Cryo-EM) generates up to tera-scale datasets per single materials [20, 22]. The scalability and efficiency to process the large volume of data are clearly important for the nationwide service deployment.
@@
The reminder of this paper is summarized as follows. Section II illustrates our governmental project, facility centers and supercomputers as the background of our work. Section III discusses the overview of the system. In Section IV, the implementation of the system is described. In Sections V and VI, we show the basic performance result and several use cases. Finally, in Sections VIIVIII and IX, we summarize our key findings, the suggested future directions and related work, before concluding this paper.
II Background
This section discusses a governmental project on materials science, experimental material facility centers, an academic cloud, and supercomputers in Japan, which are related to the ARIM-mdx data system.
II-A ARIM Japan Project
@@
Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM Japan 2021–2031) [28] is a continuation of a series of long-term initiatives (Nanotechnology Network Project in 2007–2011, Low Carbon Research Network Project in 2011–2016, Nanotech Technology Platform Project in 2012–2021) that aim to effectively manage shared-use experimental materials-science facilities located broadly in 13 universities and research institutes to support the fundamental research of material science and nanotechnology. Over 1000 units of experimental equipment are registered for shared use, and more than 2000 research projects utilize them annually. The areas of the research projects cover a wide range of topics, including but not limited to semiconductors and photonic materials, biomaterials, catalytic materials, and energy storage materials.
II-B Shared-use Materials Research Facility Centers in Japan
The ARIM-mdx data system has been involving several materials research facilities for shared use as shown in Fig. 2. We briefly summarize these facilities as follows:
1. Inst. of Eng. Innovation at The Univ. of Tokyo: In the Institute of Engineering Innovation (IEI) at the University of Tokyo [29], more than 30 units of experimental equipment on nanostructure measurement and analysis are available for shared use, including (Scanning) Transmission Electron Microscopy (STEM/TEM), X-ray photoelectron spectroscopy (XPS), Scanning electron microscope (SEM), and several physical property measurement systems. Over 600 users from 300 projects utilize the facility annually.
2. Takeda Super Clean Room at The Univ. of Tokyo: Takeda Super Clean Room [30] is one of the largest and cleanest Class 3 clean rooms located in Tokyo (Class 3: no more than one particle larger than 0.5 µm in one cubic foot of air, as defined in ISO14644-1 [31]). Over 80 units of experimental equipment mainly on nanofabrication are installed for shared use, including, lithography/exposure/drawing, film formation/deposition/etching, and so forth. Approximately 1000 users from 190 projects utilize the facility annually.
3. SPring-8: SPring-8 [32] is the world’s largest third-generation synchrotron radiation facility. Synchrotron radiation is powerful electromagnetic radiation, which is emitted when near-light-speed electrons are curved by a magnetic field. The research fields utilizing SPring-8 cover a wide variety of topics including nanotechnology, biotechnology and the various industrial applications. Over 2000 research projects are conducted annually in total, and a part of them are managed by the ARIM project.
4. Research Inst. for Semiconductor Eng. at Hiroshima Univ.: In Research Institute for Semiconductor Engineering (RISE) at Hiroshima University [33], around 50 units of equipment on nanofabrication are installed for shared uses. Around 30 research projects are conducted annually, with a main focus on silicon carbide (SiC) semiconductors and power devices.
5. Nanocharactreization Facility at AIST: In Nanocharactreization Facility at National Institute of Advanced Industrial Science and Technology (AIST) [34], more than 60 units of material experimental equipment are registered as shared use, with a wide range of equipment for nano-processing and nano-characterization.
6. Creative Research Institution at Hokkaido Univ.: In Creative Research Institution at Hokkaido University [35], over 50 units of material experimental equipments are installed for shared use. It includes various facilities for nanostructure measurement and analysis, as well as nanofabrication in a clean room.
7. Ultramicroscopy Research Center at Kyushu Univ.: In Ultramicroscopy Research Center at Kyushu University [36], more than 30 units of materials research equipment are offered as shared use, mainly on electron microscopies (TEM, STEM and SEM). The research fields that use these facilities are diverse, with a particular focus on the nanoscale materials such as nanotubes and nanosheets.
8. Inst. of Mater. and Syst. for Sustainability in Nagoya Univ.: The Institute of Materials and Systems for Sustainability (IMaSS) in Nagoya Univ. [37] offers over 80 units of shared-use experimental equipment, including those for nanostructure measurement and nanofabrication. The research fields that use these facilities are diverse, with a particular focus on the biological sciences.
9. The Coordinated Center for UEC Research Facilities: In the Coordinated Center for UEC Research Facilities at The University of Electro-Communications [38], around 20 units of material experimental equipment are available for shared use, including electron microscopes and the physical property measurement systems.
Figure 2: Material research facility centers and supercomputers integrated with ARIM-mdx data system.
II-C The mdx cloud and Supercomputers in Japan
The mdx cloud system is an academic cloud computing service in Japan [39]. It provides a high-performance virtual machine (VM) service and is connected to SINET6 [40], Japanese ultra high-speed academic network infrastructure. The network accesses from each university and national research institute in Japan go far beyond the capabilities of the public internet.
There are a number of academic supercomputers in Japan. These supercomputers are also connected via SINET6. As shown in Fig. 2, supercomputers that are actually linked with the ARIM-mdx data system include RIKEN Fugaku [41] (The Japanese 1st and the world’s 4th supercomputer in TOP 500 2024), The Univ. of Tokyo Wisteria [42], Hokkaido Univ. Grand-Chariot [43], Tokyoku Univ. MASAMUNE-IMR [44], Osaka Univ. Octpus [45], and Kyushu Univ. ITO [46].
@@ III ARIM-mdx Data System: Overview
In this section, we describe the design principles and overview of our proposed system as shown in Fig. 3.
@@ III-B System Overview
Fig. 3 shows the system overview. The ARIM-mdx data system provides (i) interactive computing services, (ii) large-volume cloud storage, (iii) direct data transfer from experimental equipment, and (iv) high-throughput networking from the supercomputers in Japan. The interactive computing services and cloud storage are provided via web-based interfaces. The specifications of the interactive computation services can be flexibly configured for mid- to high-performance environments, including multiple GPUs or many cores. Direct data transfer is enabled via data-transfer IoT devices [48], which provide an offloading method to remotely transfer data from non-networked standalone computers as well as par-user data distribution with user authentication.
@@ IV-A Cloud Internal Direct Connection
The primary system-level challenge in this architecture is access latency between the mdx cloud, where Jupyter environment is deployed for interactive data analysis, and the peta-scale dedicated data storage. Since data on the dedicated storage are interactively accessed from the Jupyter on the mdx side, very low latency access between them are necessary for usability. In a conventional networking approach, however, the connection between the mdx cloud and storage would typically route through an external network gateway. This routing would result in inter-cluster latency, even if the mdx cloud and storage systems are physically located in proximity to each other.
To solve the issue, the ARIM-mdx data system establishes a cloud internal direct connection, which creates a direct, dedicated physical network between the mdx cloud and the dedicated storage. By utilizing layer-2 network virtualization of mdx, we extend the virtual cluster on the mdx to one of its physical ports. Then, by connecting this port to the dedicated storage, we establish the virtualized intra-network without any affection to the other virtual clusters in the mdx cloud. The key idea is similar to AWS Direct Connect, Google Dedicated Network, and Azure ExpressRoute, but differs in that it connects directly to the mdx’s internal network by bypassing the external network gateway.
@@ IV-B IoT-based Networking of Standalone Material Facilities
In the ARIM-mdx data system, we widely install data transfer IoT devices [48] for the direct data transfer from such standalone facility computers. The key idea behind the networking is to utilize USB volume interface of the standalone facility computers. The IoT device mimics the USB flash drive to the facility side (using USB OTG g_mass_gadget) while these are remotely connected to the cloud storage. Thus, from user side on facility operation computers, the IoT devices are recognized as USB flash memory, where users just copy their experimental data to the fake USB volume folder. Once the data are copy to this fake USB volume folder, the IoT devices automatically transfer data to the cloud storage of the system. Furthermore, by integrating this USB-based data transfer with user authentication, the system can remotely and safely distribute the experimental data to the storage in per-user manner, without any (Ethernet-based) internet connection in the facility computers.
@@ V System Evaluation and Statistic
In this section, we briefly evaluate the performance of the key technological functions in the ARIM-mdx data system. Then, statistical information for this one-year trial operation is shown.
V-A Evaluation for Cloud Internal Direct Connection and Wide-Area Networking
@@
  • The system achieves very low access latency (less than 1ms) from Jupyter service to the storage system, which is 4.7 times faster than conventional approaches and provides latency performance on par with intra-cluster environments.
  • The network connection from the experimental facilities and supercomputers to the ARIM-mdx data system achieves 4.0 times higher throughput than public cloud services.
The mdx cloud system includes 100 Gb/s networks for both the external and internal connections. SINET6 is built on a 400Gb/s network infrastructure. The data storage system runs Nextcloud 29.0 powered by Red Hat Enterprise Linux release 8.5 with CPU Intel(R) Xeon(R) Gold 6348 $\times$ 2 (28 cores $\times$ 2 Sockets) and 64 GB Memory. Approximately 3 petabyte DDN EXAScaler Storage (Lustre) are installed for the storage system. The physical network connection between the mdx cloud and the storage system is based on 10GbE network switches.
Fig. 6 shows the ping access latency from various sources to the storage system of ARIM-mdx data system. We conducted the test five times and report the median value. The variation between tests was less than 5%. ARIM-mdx Jupyter shows the access latency from the Jupyter environment on the mdx cloud to the dedicated storage system. Its value is 0.87 milliseconds (ms). Thanks to our established cloud internal direct connection, its access latency is the similar performance to the intra-cluster environment (e.g., typically 0.30.5 ms). Inside-Campus Server shows the access latency from the physically located in the same building as the mdx cloud (but the network is not directly connected). Its value is 4.24 ms due to the fact that the network is through the general route to the mdx cloud. The packet is once routed to the regional upstream and then goes down to the mdx cloud through the external network gateway. Wisteria (East Area) shows the access latency from Wisteria supercomputer in The University of Tokyo. Its value is 4.13 ms due to the fact that Wisteria is located in the same building as the mdx cloud. Fugaku (West Area) shows the latency from Fugaku supercomputer, which is located in Kobe, the west area of Japan. Due to the physical distance between Kobe and Tokyo, its value becomes 11.9 ms. MS Azure Japan East and MS Azure Japan West show the latency from virtual machines on Japan East region and Japan West region in Mincrosoft Azure service. As the storage server is located in Tokyo area (the east side of Japan), the latencies become 4.85 ms for Japan East and 12.03 ms for Japan West, respectively.
@@
Fig. 6 shows the data transfer throughput between each server and the storage system of the ARIM-mdx data system. We transferred 10 files of 1GB each (10GB total) five times using Rclone copy (ver. 1.67 with 4 parallel transfers) and report the median value, with errors less than 5ARIM-mdx Jupyter shows the throughput between Jupyter container and the storage system on ARIM-mdx data system. Thanks to the cloud direct connection, its throughput shows the best performance. Its value is 598.8 megabyte per second (MByte per sec). Inside-Campus Server shows the worst performance due to the bottleneck at its commodity network switches connecting to the campus network. Even though the intra-campus network uses high-performance infrastructure, the throughput remains low. Its value is 51.65 MByte per sec. Wisteria (East Area) shows the performance between Wisteria supercomputer and ARIM-mdx data system (storage server). The connection from Wisteria to ARIM-mdx data system uses high-performance infrastructure including SINET6. Thus, it achieves high-throughput and its value is 425.5 MByte per sec. Similar to Wisteria, Fugaku (West Area) also achieves very high throughput due to the SINET6. Its value is 512.8 MByte per sec. The slight performance difference between Wisteria and Fugaku may be due to the specification of their storage systems and login nodes where the program was executed. MS Azure Japan East and MS Azure Japan West show the throughput from the VM machines in Microsoft Azure, where we use Standard_D4_v3 instance type (4 vCPUs and 16 memory). Their values are 128.3 and 128.0 MByte per sec, respectively. Their lower throughput is due to the public Internet and the bandwidth restriction imposed by the cloud provider.

### v2

@@ Abstract
In modern materials science, effective and high-volume data management across leading-edge experimental facilities and world-class supercomputers is indispensable for cutting-edge research. However, existing integrated systems that handle data from these resources have primarily focused just on smaller-scale cross-institutional or single-domain operations. As a result, they often lack the scalability, efficiency, agility, and interdisciplinarity, needed for handling substantial volumes of data from various researchers.
In this paper, we introduce ARIM-mdx data system [^1] The service is available at https://arim.mdx.jp., aiming at a nationwide data platform for materials science in Japan. Currently in its trial phase, the platform has been involving 11 universities and institutes all over Japan, and it is utilized by over 800 researchers from around 140 organizations in academia and industry, being intended to gradually expand its reach. The ARIM-mdx data system, as a pioneering nationwide data platform, has the potential to contribute to the creation of new research communities and accelerate innovations.
@@
Big Data, Materials Science, Supercomputer, Cloud Computing, Shared-use Research Facility
@@ I Introduction
Figure 1: Collaboration between experimental and theoretical materials science. The example shows a typical nanostructure analysis, where the atomic-scale image of the targeting material is obtained from the electron microscope whereas it is analyzed via computation-intensive theoretical simulation using the supercomputer to detect its structure. The raw image is obtained from [2].
@@
However, despite the substantial success in the wide collection and large-scale management of the computational data, that of the real experimental data remains a significant challenge. Most efforts on the collection, management and analysis of the experimental data are primarily focused just on the smaller-scale cross-institution [13, 14, 15, 16, 17, 18, 19, 20], single domain [21, 22, 23, 24, 25, 26], or data curated from research articles [27, 28]. This leaves a huge potential for leveraging a large amount of the raw, unprocessed data untouched.
To address these gaps, we have been developing ARIM-mdx data system, aiming at a nationwide materials-science data infrastructure capable of effectively handling both computational and experimental data. The ARIM-mdx data system is designed to meet essential system requirements for data storage, management, and analysis. The system requirements and our contributions towards this nationwide deployment are summarized as follows:
@@
For the stability and security, the ARIM-mdx data system is based on a peta-scale dedicated storage system, rather than on public clouds or shared-use storage systems. Such an isolated setup enables us to make physical operations be fully under control. Moreover, to ensure safe data transfer from network-disabled facility computers, we offload the data transfer mechanism to IoT devices, thereby mitigating security risks on the facility computers.
Scalability and Efficiency: Due to the rapid advancement of experimental facilities and supercomputers, these generate richer data at a higher rate. For example, the state-of-the-art Cryogenic Electron Microscopy (Cryo-EM) generates up to TB-scale datasets per single material [21, 23]. The scalability and efficiency to process the large volume of data are clearly important for the nationwide service deployment.
@@
The remainder of this paper is summarized as follows. §II illustrates the background. §III discusses the overview of the system. In §IV, the implementation of the system is described. In §V and VI, we show the basic performance result and several use cases. Finally, in §VIIVIII and IX, we summarize our key findings, the suggested future directions and related work, before concluding this paper.
II Background: ARIM Japan Project
@@
Advanced Research Infrastructure for Materials and Nanotechnology in Japan (ARIM Japan 2021–2031) [29] is a continuation of a series of long-term initiatives (Nanotechnology Network Project in 2007–2011, Low Carbon Research Network Project in 2011–2016, Nanotech Technology Platform Project in 2012–2021), that aim to effectively manage shared-use experimental facilities located broadly in 13 universities and research institutes. Over 1,000 units of experimental equipment are registered for shared use, and more than 2,000 research projects utilize them annually.
@@ III-B System Overview
Fig. 2 shows the system overview. The ARIM-mdx data system provides (i) interactive computing services, (ii) large-volume cloud storage, (iii) direct data transfer from experimental equipment, and (iv) high-throughput networking from the supercomputers in Japan. The interactive computing services and cloud storage are provided via web-based interfaces. The specifications of the interactive computation services can be flexibly configured for mid- to high-performance environments, including multiple GPUs or many cores. Direct data transfer is enabled via data-transfer IoT devices [31], which provide an offloading method to remotely transfer data from non-networked standalone computers as well as per-user data distribution with user authentication.
@@ III-B3 As a service for managing and analyzing supercomputer-generated data
III-C Universities and Institutions Involved
The ARIM-mdx data system has been involving several universities and institutes which operate research facilities and/or supercomputers as follows: (1) the University of Tokyo, (2) SPring-8: the world’s largest third-generation synchrotron radiation facility, (3) Hiroshima University, (4) RIKEN Fugaku: the Japanese 1st and the world’s 4th supercomputer in TOP 500 2024, (5) Hokkaido University, (6) Tohoku University, (7) Science Tokyo, (8) Nagoya University, (9) Kyoto University, (10) Osaka University, and (11) Kyushu University.
@@ IV-A Cloud Internal Direct Connection
The primary system-level challenge in this architecture is access latency between the mdx cloud, where the Jupyter environment is deployed for interactive data analysis, and the peta-scale dedicated data storage. Since data on the dedicated storage are interactively accessed from the Jupyter on the mdx side, very low latency accesses between them are necessary for usability. In a conventional networking approach, however, the connection between the mdx cloud and storage would typically route through an external network gateway. This routing would result in inter-cluster latency, even if the mdx cloud and storage systems are physically located in proximity to each other.
To solve the issue, the ARIM-mdx data system establishes a cloud internal direct connection, which creates a direct, dedicated physical network between the mdx cloud and the dedicated storage. By utilizing layer-2 network virtualization of mdx, we extend the virtual cluster on the mdx to one of its physical ports. Then, by connecting this port to the dedicated storage, we establish the virtualized intra-network without any impact on the other virtual clusters in the mdx cloud. The key idea is similar to AWS Direct Connect, Google Dedicated Network, and Azure ExpressRoute, but differs in that it connects directly to the mdx’s internal network by bypassing the external network gateway.
@@ IV-B IoT-based Networking of Standalone Material Facilities
In the ARIM-mdx data system, we have widely installed data transfer IoT devices [31] for the direct data transfer from such standalone facility computers. The key idea behind the networking is to utilize USB volume interface of the standalone facility computers. The IoT device mimics the USB flash drive to the facility side (using USB OTG g_mass_gadget) while the devices are remotely connected to the cloud storage. Thus, from user’s side on facility operation computers, the IoT devices are recognized as USB flash memory, where users just copy their experimental data to the fake USB volume folder. Once the data are copied to this fake USB volume folder, the IoT devices automatically transfer data to the cloud storage of the system. Furthermore, by integrating this USB-based data transfer with user authentication, the system can remotely and safely distribute the experimental data to the storage in per-user manner, without any (Ethernet-based) internet connection in the facility computers.
@@ V System Evaluation and Statistic
V-A Evaluation for Key Components
@@
  • The system achieves very low access latency (less than 1ms) from Jupyter service to the storage system, which is 4.7 times faster than conventional approaches.
  • The network connection from the experimental facilities and supercomputers achieves 4.0 times higher throughput than public cloud services.
The mdx cloud system includes 100 Gb/s networks for both the external and internal connections. SINET6 is built on a 400Gb/s network infrastructure. The data storage system runs Nextcloud 29.0 powered by Red Hat Enterprise Linux release 8.5 with CPU Intel(R) Xeon(R) Gold 6348 $\times$ 2 (28 cores $\times$ 2 Sockets) and 64 GB Memory. Approximately 3 petabytes DDN EXAScaler Storage (Lustre) is installed. The physical network connection between the mdx cloud and the storage system is based on 10GbE network switches.
@@
Fig. 5 shows the ping access latency from various sources to the storage system of ARIM-mdx data system. We conducted the test five times and report the median value. The variation between tests was less than 5%. ARIM-mdx Jupyter shows the access latency from the Jupyter environment on the mdx cloud to the dedicated storage system. Its value is 0.87 milliseconds (ms). Thanks to our established cloud internal direct connection, its access latency is similar performance to the intra-cluster environment (e.g., typically 0.30.5 ms). Inside-Campus Server shows the access latency from the server, which is physically located in the same building as the mdx cloud (but the network is not directly connected). Its value is 4.24 ms due to the fact that the network is through the general route to the mdx cloud. The packet is once routed to the regional upstream and then goes down to the mdx cloud through the external network gateway. Wisteria (East Area) shows the access latency from Wisteria supercomputer in the University of Tokyo. Its value is 4.13 ms due to the fact that Wisteria is located in the same building as the mdx cloud. Fugaku (West Area) shows the latency from Fugaku supercomputer, which is located in Kobe, the western area of Japan. Due to the physical distance between Kobe and Tokyo, its value becomes 11.9 ms. MS Azure Japan East and MS Azure Japan West show the latency from virtual machines on the Japan East region and Japan West region in Microsoft Azure service. As the storage server is located in the Tokyo area (the east side of Japan), the latencies become 4.85 ms for Japan East and 12.03 ms for Japan West, respectively.
Fig. 5 shows the data transfer throughput between each server and the storage system of the ARIM-mdx data system. We transferred 10 files of 1GB each (10GB total) five times using Rclone copy (ver. 1.67 with 4 parallel transfers) and report the median value, with errors less than 5%. ARIM-mdx Jupyter shows the throughput between Jupyter container and the storage system on ARIM-mdx data system. Thanks to the cloud direct connection, its throughput shows the best performance. Its value is 598.8 megabyte per second (MByte per sec). Inside-Campus Server shows the worst performance due to the bottleneck at its commodity network switches connecting to the campus network. Even though the intra-campus network uses high-performance infrastructure, the throughput remains low. Its value is 51.65 MByte per sec. Wisteria (East Area) shows the performance between Wisteria supercomputer and ARIM-mdx data system (storage server). The connection from Wisteria to ARIM-mdx data system uses high-performance infrastructure including SINET6. Thus, it achieves high-throughput and its value is 425.5 MByte per sec. Similar to Wisteria, Fugaku (West Area) also achieves very high throughput due to the SINET6. Its value is 512.8 MByte per sec. The slight performance difference between Wisteria and Fugaku may be due to the specification of their storage systems and login nodes where the program was executed. MS Azure Japan East and MS Azure Japan West show the throughput from the VM in Microsoft Azure, where we use Standard_D4_v3 instance type (4 vCPUs and 16 GB memory). Their values are 128.3 and 128.0 MByte per sec, respectively. Their lower throughput is due to the public Internet and the bandwidth restriction imposed by the cloud provider.